# ACoRN: Noise-Robust Abstractive Compression in Retrieval-Augmented Language Models

### v1

@@ Abstract
Index Terms:
Noise robustness, Abstractive compression, Retrieval-augmented language models, Large language models
@@ V-A Main Results
Figure 4: Comparison of GPT-3.5-turbo QFS performance when only evidential documents are included in the prompt versus when all top-5 documents are included, based on random sampling of 100 cases for each evidential document count $N$ in top-5 retrieval. When $N$=0 with retrieved only evidential documents means using only internal knowledge. The compressed output is passed to the inference model’s prompt, with the language model $M$ being LLaMA-3.1-8B-Instruct. The dotted line represents the performance when summarization is done by randomly sampling 100 instances, regardless of $N$.
@@ V-B Reliability of Summarization Labels
As shown in Fig. 4, we compared the reliability of summarization labels generated using all documents versus using only evidential documents. We split the test dataset based on the number of evidential documents $N$ from the top-$5$ retrieved documents. We aim to examine how the quality of summarization labels changes with the proportion of presented evidential documents and how the performance gap between the two settings varies accordingly. We randomly sampled 100 instances for each $N$ to ensure fair evaluation. We observe the summarization performance of the teacher model in two scenarios: when only $N$ evidential documents are provided, and when $N$ evidential documents are combined with $5-N$ noise documents. When the fixed value of $N$ is small, using only the evidential documents for summarization is more effective compared to using all retrieved documents. When we increase $N$, there are cases where utilizing all retrieved documents is more effective. This is because the retrieval noise from documents with low semantic relevance to the query can help better differentiate evidential documents from noise, leading to a more effective summarization [18]. However, the effect is minimal when compared to using only evidential documents. Results show that using only evidential documents for summarization labels is more effective.

### v2

@@ V-A Main Results
Figure 4: Comparison of GPT-3.5-turbo QFS performance when only evidential documents are included in the prompt versus when all top-5 documents are included, based on random sampling of 100 cases for each evidential document count NN in top-5 retrieval. When NN=0 with retrieved only evidential documents means using only internal knowledge. The compressed output is passed to the inference model’s prompt, with the language model $M$ being LLaMA-3.1-8B-Instruct. The dotted line represents the performance when summarization is done by randomly sampling 100 instances, regardless of NN.
@@ V-B Reliability of Summarization Labels
As shown in Fig. 4, we compared the reliability of summarization labels generated using all documents versus using only evidential documents. We split the test dataset based on the number of evidential documents $N$ from the top-$5$ retrieved documents. We aim to examine how the quality of summarization labels changes with the proportion of presented evidential documents and how the performance gap between the two settings varies accordingly. We randomly sampled 100 instances for each $N$ to ensure fair evaluation. We observe the summarization performance of the teacher model in two scenarios: when only $N$ evidential documents are provided, and when $N$ evidential documents are combined with 5N5-N noise documents. When the fixed value of $N$ is small, using only the evidential documents for summarization is more effective compared to using all retrieved documents. When we increase $N$, there are cases where utilizing all retrieved documents is more effective. This is because the retrieval noise from documents with low semantic relevance to the query can help better differentiate evidential documents from noise, leading to a more effective summarization [18]. However, the effect is minimal when compared to using only evidential documents. Results show that using only evidential documents for summarization labels is more effective.